# Non-continuous valuations on convex bodies and a new characterization of volume

### v1

@@ Abstract.
Research partially supported by grants PID2020-116398GB-I00, PID2024-162214NB-I00 and CEX2023-001347-S funded by MCIN/AEI/10.13039/501100011033.
@@ 1. Introduction
The theory of valuations on convex bodies has become one of the central pillars of modern convex and integral geometry. Its origins can be traced back to the early 20th century, when Dehn proved a negative solution for Hilbert’s Third Problem [Dehn]. Early contributions by Hadwiger, most notably his characterization theorem in the 1950s, established a cornerstone of the field: every continuous, rigid motion invariant valuation on convex bodies in Euclidean space is a linear combination of intrinsic volumes (see [Hadwiger]). This result revealed the deep structure underlying seemingly disparate geometric measures and opened the door to systematic classification results.
In the decades that followed, the theory of valuations underwent a remarkable expansion, influenced by methods from topology, measure theory, and representation theory. The introduction of continuous and smooth valuations by Alesker in the early 2000s [Alesker1] marked a new phase in the subject, connecting convex geometry with areas such as integral geometry on manifolds and the theory of distributions. As a result, valuation theory has become an active interface between convex geometry, geometric measure theory, and algebraic topology, with applications ranging from stochastic geometry to complex and Hermitian integral geometry (see, e.g. [Bernig, Handbook, equivariantes, LibroHug, MonikayFabian, logconcavas, Schneider]). In the recent years, relevant classification results on valuation theory on different context have appeared (see [Alesker1999, BernigFu2011, ColesantiLudwigMussnig2024, Haberl, Knoerr24, LudwigReitzner1999, LudwigReitzner2010, Estrellado2, Reticulos]).
@@
In this direction, it is worth noting the theory of automatic continuity, which emerged from the study of algebraic structures endowed with compatible topologies, where one seeks to understand when a purely algebraic property together with appropriate definability conditions actually suffice to guarantee continuity of certain maps. In its most classical form, the theory addresses the question of when a homomorphism between topological groups (see [RosendalAC]), rings, or algebras must necessarily be continuous, even without explicit continuity assumptions. These foundational insights reveal that, in many analytical contexts, continuity is not an additional hypothesis but rather an inevitable consequence of the underlying structure.
The theory of valuations and the theory of automatic continuity share a foundational element in their early development: The Cauchy Functional Equation. In the former, it underlies the basis for the definition of what is considered the first valuation, Dehn’s valuation [Dehn]. In the latter, it is the primordial example of the theory and the first example studied in depth [Cauchy], [Hamel]. Although both theories share this important link, to our knowledge not much attention has been paid to the use of automatic continuity in the theory of valuations.
In this article, we investigate the applicability of techniques from automatic continuity theory to valuation theory, with the aim of replacing the continuity condition with weaker conditions, such as boundedness or Borel measurability, that might still guarantee some control on the behavior of valuations and allow us to find useful characterizations.
@@ 2.1. Preliminaries on valuations
where $B$ denotes the closed Euclidean unit ball. We refer to [Schneider] as a standard reference on these notions. Let us denote by $\mathcal{K}(B)$ the set of convex bodies lying inside the unit closed unit ball $B$, that is
@@
Most of the valuations considered in this paper will be scalar-valued. However, we will also make use of area measures, denoted by $S_{i}$, which are valuations taking values on $\mathcal{M}(\mathbb{S}^{n-1})$, the space of Borel regular measures on the sphere. For a detailed account of these valuations, we refer to [Schneider, Chapter 4].
@@
When working in $\mathcal{P}^{n}$ or in $\text{Pa}(\{e_{i}\}_{i=1}^{n})$, it is known that weak additivity is equivalent to the valuation property (see [Schneider, Theorem 6.2.3]). This is no longer the case for $\mathcal{K}^{n}$ (see [Debiladitiva, Page 173]); however, if $V$ is weakly additive on $\mathcal{K}^{n}$, then its restrictions, $V|_{\mathcal{P}^{n}}$ and $V|_{\text{Pa}(\{e_{i}\}_{i=1}^{n})}$, are valuations on their respective domains. This fact allows us to weaken the valuation hypothesis to the condition that $V$ is weakly additive.
@@
It is well known that all valuations defined on $\mathcal{P}^{n}$ satisfy the Inclusion-Exclusion property (see [Schneider, Theorem 6.2.3]). In contrast, in the case of $\mathcal{K}^{n}$, it is known that all continuous valuations satisfy this (in fact, $\sigma$-continuity is enough, see [sigmacont, Theorem 1.21]). However, it remains an open question whether continuity can be completely dropped. Since our focus will be on valuations which are not necessarily continuous, we will not assume Inclusion-Exclusion.
Let us recall the fundamental volume characterization theorem for valuations due to Hadwiger [Hadwiger].
@@
Our main result (Theorem 1.1) is an extension of this theorem where we replace the continuity hypothesis by a (weaker) boundedness condition. We refer the reader to [Hadwiger] for the original proof of Theorem 2.5 and to [Alesker, Theorem 3.1.1] for a classical proof using valuation theory. The original proof in [Hadwiger] is based in the theory of contents while the proof in [Alesker] is based in McMullen’s decomposition theorem, which we recall next:
@@
Motivated by this, we explore whether the continuity assumptions can be relaxed in other classical results from valuation theory. Recall McMullen’s characterization of $\mathbf{Val}_{n-1}$ (see [McMullen]):
@@ 2.2. Preliminaries on automatic continuity
for any $x,y\in\mathbb{R}$. Solutions to this equation are called additive functions. Clearly, all linear functions satisfy the Cauchy Functional Equation, but could these be the only solutions? Cauchy [Cauchy] proved that if the function $f$ is continuous, then the answer is positive. Later, in 1906, Hamel [Hamel] proved the existence of discontinuous additive functions by constructing a basis of $\mathbb{R}$ over $\mathbb{Q}$, now known as a Hamel Basis. These discontinuous additive functions are also known as Hamel functions. An important aspect of the Cauchy Functional Equation is the dichotomy in the nature of its solutions, either they are continuous and linear, or extremely pathological functions. This fact can be summarized in the following well-known theorem (see e.g. [Kuczma]).
@@
A very complete survey of automatic continuity on the more general setting of group homomorphisms can be found in [RosendalAC] (see also [RosendalPI] for more recent developments). Theorems in the spirit of 2.8 is what we refer to as an automatic continuity result, and although most developments usually focus on the Borel condition, our interest in the setting of valuations will shift to the boundedness condition.
@@ 3. Translation Invariant valuations on parallelotopes
In this section, we first provide a characterization of translation invariant valuations on $\text{Pa}(\{e_{i}\}_{i=1}^{n})$ without assuming any continuity. Although the proof can be considered elementary, we have not found any explicit reference. This characterization will prove useful and it will show that translation invariant valuations do satisfy some analog to Theorem 2.8 (see Theorem 3.9). We refer to [Klain, Ch. 4] for a study of continuous and translation invariant valuations on $\text{Pa}(\{e_{i}\}_{i=1}^{n})$.
@@
Theorem 3.6 can be interpreted as a version of McMullen’s Theorem on decomposition for valuations on polytopes (see [Alesker]). We characterize the homogeneous parts of the decomposition as separately additive functions of different dimensions, specifically when working within $\text{Pa}(\{e_{i}\}_{i=1}^{n})$. This formulation serves as a substitute for McMullen’s Theorem and we will use Theorem 3.6 in the proof of Theorem 1.1. This also leads to the following automatic continuity result.
@@
where we used the fact that additive functions are $\mathbb{Q}$-linear [Kuczma, Theorem 5.2.1]
@@ 4. A new characterization of the volume in $\mathbb{R}^{n}$
where $S(K)$ denotes the surface area of $K$ (see [Schneider, Chapter 4]). On the other hand, since $B(0,r)\subset K\subset B(0,R)$, for $\varepsilon$ small enough we get
@@
(see [Schneider, Lemma 2.3.6]). Combining these results we get
@@
Again, we restate that in the previous Lemma 4.3 the valuation condition can be relaxed, we only need the mapping $V$ to be weakly additive. This property is strictly weaker than being a valuation (see [Debiladitiva, Page 173] for an example of a weakly additive function on $\mathcal{K}^{n}$ which is not a valuation). Nevertheless, if $V$ is weakly additive, it is straightforward to see that the restriction $V|_{\text{Pa}(\{e_{i}\}_{i=1}^{n})}$ is a valuation on $\text{Pa}(\{e_{i}\}_{i=1}^{n})$, allowing us to apply the results from Section 3
@@ 5. Degrees of homogeneity for a bounded valuation
To see that this set can not be reduced any further, we construct bounded and translation invariant valuations, $\phi_{l}$, that are $m$-homogeneous for any $m\in[0,n-1]$. These examples show both that each degree in this interval actually occurs, and that the argument from Lemma 4.3 can not be pushed to extend the values of $m$. Indeed, each $\phi_{l}$ vanishes on polytopes while satisfying $\phi_{l}(B)>0$. These valuations are analogous to the $p$-affine surface areas (see [MonikayFabian, Section 3.3] for a gentle introduction) and are based on the Gauss Kronecker curvature.
@@
A well known result due to Aleksandrov (see [Alexandroff] or [Schneider, Theorem 2.6.1]) states that the boundary of a convex body is twice differentiable almost everywhere with respect to the (n1)(n-1)-dimensional Hausdorff measure. This implies that the curvature is well defined almost everywhere, and moreover, $x\mapsto\kappa(K,x)$ is measurable. This mapping can also be described as the density of the absolutely continuous part of the curvature measure $C_{0}(K,\cdot)$ with respect to the (n1)(n-1)-dimensional Hausdorff measure. We refer the reader to [Schneider, Chapter 4] and [Hug, Theorem 3.2] for the definition of $C_{0}(K,\cdot)$ and details. As a consequence, we have the following.
@@
where $\chi_{A}(x)$ denotes the indicator function of the set $A$. Applying the property [pushforward, Theorem 3.6.1] we get that
@@
This is well defined since $\kappa(K,\cdot)$ is measurable and nonnegative, also, in the case $l=0$, we adopt the convention that $0^{0}=0$. We will need the following lemma taken directly from [Schutt, Lemma 5].
@@
With this lemma and following the same procedure as in [Schutt] we can prove that the previous mapping is a valuation.
@@ 6. Some negative results
The proof is based on [McMullen, Theorem 1], with minor modifications, we include it here for completeness.
@@
Let $u=\sum_{i=1}^{n}\eta_{i}e_{i}\in\mathbb{S}^{n-1}$ with $\eta_{i}\neq 0$ for all $i=1,\dots,n$. We take $e_{i}^{\prime}=-\mathrm{sign}(\eta_{i})e_{i}$ and a simplex $T$ with $u,e_{1}^{\prime},\dots,e_{n}^{\prime}$ as its normal vectors. By [Schneider, Lemma 5.1.1], taking $F_{i}=F(T,e_{i}^{\prime})$ and $F_{0}=F(T,u)$, we have
@@
Similarly, the mapping $\mu\mapsto\mu^{s}$, which gives the singular part of any measure with respect to the Haar measure, is linear. This follows from the uniqueness of the Lebesgue decomposition [Cohn, Theorem 4.3.2]. Finally, the last mapping is defined as
@@
Besides Minkowski addition, Blaschke addition is another semigroup operation inside $\mathcal{K}^{n}$, that has proven to be quite useful. For the definition we refer the reader to [Schneider] and [McMullen]. We will see that a valuation that is Blaschke additive and bounded does not have to be continuous. To exhibit such an example, we will make use of McMullen’s Characterization of (n1)(n-1)-homogeneous continuous and translation invariant valuations, Theorem 2.7. The idea is to build valuations via the surface area measure, $S_{n-1}$, which is Blaschke additive, and then modify the function in the integral representation to obtain a valuation that is Borel measurable and bounded but not continuous.
@@
for any $K\in\mathcal{K}^{n}$. Since any balanced measure can be written as the difference of two surface area measures (check [McMullen, Theorem 3]) we get that
@@
Now, let $L=\{\langle u,\cdot\rangle,u\in\mathbb{R}^{n}\}\subset C(\mathbb{S}^{n-1})$ be the subspace of linear functions. Note that this has finite dimension, which means that it is closed in all vector topologies. Now, the polar of $L$ is exactly the set of balanced measures and fgf-g is in the polar of balanced measures. By the Bipolar Theorem (see [Fabian, Theorem 3.38]) we get that fgf-g is in $L$, that is, fgf-g is continuous and, since $g$ was continuous, we get that $f$ is continuous.
The converse result is immediate from the fact (see [Schneider, Theorem 4.2.1]) that convergence of convex bodies $K_{j}\xrightarrow[]{}K$ in the Hausdorff metric implies that, for every continuous $f:\mathbb{S}^{n-1}\to\mathbb{R}$,
@@
Therefore, $\phi_{\chi_{\bigcup_{j\in\mathbb{N}}A_{j}}}$ is the pointwise limit of measurable functions and thus measurable. These facts show that $\mathcal{D}$ is a Dynkin class that contains all open sets. By a classical result of Dynkin classes (check [Cohn, Chapter 1.6]), all Borel sets belong to $\mathcal{D}$.

### v2

@@ Abstract.
Key words and phrases:
Bounded valuation; Homogeneous valuation; Convex bodies; Volume
2020 Mathematics Subject Classification:
52B45,52A38
Research partially supported by grants PID2020-116398GB-I00, PID2024-162214NB-I00 and CEX2023-001347-S funded by MCIN/AEI/10.13039/501100011033. The first author has also received financial support from Ministerio de Ciencia, Innovación y Universidades through an FPU Grant FPU22/02055
@@ 1. Introduction
The theory of valuations on convex bodies has become one of the central pillars of modern convex and integral geometry. Its origins can be traced back to the early 20th century, when Dehn proved a negative solution for Hilbert’s Third Problem [11]. Early contributions by Hadwiger, most notably his characterization theorem in the 1950s, established a cornerstone of the field: every continuous, rigid motion invariant valuation on convex bodies in Euclidean space is a linear combination of intrinsic volumes (see [17]). This result revealed the deep structure underlying seemingly disparate geometric measures and opened the door to systematic classification results.
In the decades that followed, the theory of valuations underwent a remarkable expansion, influenced by methods from topology, measure theory, and representation theory. The introduction of continuous and smooth valuations by Alesker in the early 2000s [2] marked a new phase in the subject, connecting convex geometry with areas such as integral geometry on manifolds and the theory of distributions. As a result, valuation theory has become an active interface between convex geometry, geometric measure theory, and algebraic topology, with applications ranging from stochastic geometry to complex and Hermitian integral geometry (see, e.g. [6, 13, 19, 20, 28, 37, 40]). In the recent years, relevant classification results on valuation theory on different context have appeared (see [1, 5, 10, 16, 25, 29, 30, 42, 43]).
@@
Previous related research includes several works where non continuous valuations on polytopes are considered (see, for example, [27, 34, 31]), and other works where the focus is on non continuous valuations defined on the convex bodies which have the origin as an interior point. In this case, additional symmetry assumptions such as SL(n)(n)-invariance [14, 30] or SL(n)(n)-covariance [15] are imposed. To the best of our knowledge, non continuous translation invaluations on the full space of convex bodies, $\mathcal{K}^{n}$, have not been previously studied.
In the direction of weakening continuity conditions, it is worth recalling the theory of automatic continuity, which emerged from the study of algebraic structures endowed with compatible topologies, where one seeks to understand when a purely algebraic property together with appropriate definability conditions actually suffice to guarantee continuity of certain maps. In its most classical form, the theory addresses the question of when a homomorphism between topological groups (see [38]), rings, or algebras must necessarily be continuous, even without explicit continuity assumptions. These foundational insights reveal that, in many analytical contexts, continuity is not an additional hypothesis but rather an inevitable consequence of the underlying structure.
The theory of valuations and the theory of automatic continuity share a foundational element in their early development: The Cauchy Functional Equation. In the former, it underlies the basis for the definition of what is considered the first valuation, Dehn’s valuation [11]. In the latter, it is the primordial example of the theory and the first example studied in depth [8], [18]. Although both theories share this important link, to the best of our knowledge not much attention has been paid to the use of automatic continuity in the theory of valuations.
In this article, we investigate the applicability of techniques from automatic continuity theory to valuation theory, with the aim of replacing the continuity condition with weaker conditions, such as boundedness or Borel measurability, which might still guarantee some control on the behavior of valuations and allow us to find useful characterizations.
@@ 2.1. Preliminaries on valuations
where $B$ denotes the closed Euclidean unit ball. We refer to [40] as a standard reference on these notions. Let us denote by $\mathcal{K}(B)$ the set of convex bodies lying inside the unit closed unit ball $B$, that is
@@
Most of the valuations considered in this paper will be scalar-valued. However, we will also make use of area measures, denoted by $S_{i}$, which are valuations taking values on $\mathcal{M}(\mathbb{S}^{n-1})$, the space of Borel regular measures on the sphere. For a detailed account of these valuations, we refer to [40, Chapter 4].
@@
When working in $\mathcal{P}^{n}$ or in $\text{Pa}(\{e_{i}\}_{i=1}^{n})$, it is known that weak additivity is equivalent to the valuation property (see [40, Theorem 6.2.3]). This is no longer the case for $\mathcal{K}^{n}$ (see [35, Page 173]); however, if $V$ is weakly additive on $\mathcal{K}^{n}$, then its restrictions, $V|_{\mathcal{P}^{n}}$ and $V|_{\text{Pa}(\{e_{i}\}_{i=1}^{n})}$, are valuations on their respective domains. This fact allows us to weaken the valuation hypothesis to the condition that $V$ is weakly additive.
@@
It is well known that all valuations defined on $\mathcal{P}^{n}$ satisfy the Inclusion-Exclusion property (see [40, Theorem 6.2.3]). In contrast, in the case of $\mathcal{K}^{n}$, it is known that all continuous valuations satisfy this (in fact, $\sigma$-continuity is enough, see [22, Theorem 1.21]). However, it remains an open question whether continuity can be completely dropped. Since our focus will be on valuations which are not necessarily continuous, we will not assume Inclusion-Exclusion.
Let us recall the fundamental volume characterization theorem for valuations due to Hadwiger [17].
@@
Our main result (Theorem 1.1) is an extension of this theorem where we replace the continuity hypothesis by a (weaker) boundedness condition. We refer the reader to [17] for the original proof of Theorem 2.5 and to [3, Theorem 3.1.1] for a classical proof using valuation theory. The original proof in [17] is based in the theory of contents while the proof in [3] is based in McMullen’s decomposition theorem, which we recall next:
@@
Motivated by this, we explore whether the continuity assumptions can be relaxed in other classical results from valuation theory. Recall McMullen’s characterization of $\mathbf{Val}_{n-1}$ (see [36]):
@@ 2.2. Preliminaries on automatic continuity
for any $x,y\in\mathbb{R}$. Solutions to this equation are called additive functions. Clearly, all linear functions satisfy the Cauchy Functional Equation, but could these be the only solutions? Cauchy [8] proved that if the function $f$ is continuous, then the answer is positive. Later, in 1906, Hamel [18] proved the existence of discontinuous additive functions by constructing a basis of $\mathbb{R}$ over $\mathbb{Q}$, now known as a Hamel Basis. These discontinuous additive functions are also known as Hamel functions. An important aspect of the Cauchy Functional Equation is the dichotomy in the nature of its solutions, either they are continuous and linear, or extremely pathological functions. This fact can be summarized in the following well-known theorem (see e.g. [26]).
@@
A very complete survey of automatic continuity on the more general setting of group homomorphisms can be found in [38] (see also [39] for more recent developments). Theorems in the spirit of 2.8 is what we refer to as an automatic continuity result, and although most developments usually focus on the Borel condition, our interest in the setting of valuations will shift to the boundedness condition.
@@ 3. Translation Invariant valuations on parallelotopes
In this section, we first provide a characterization of translation invariant valuations on $\text{Pa}(\{e_{i}\}_{i=1}^{n})$ without assuming any continuity. Although the proof can be considered elementary, we have not found any explicit reference. This characterization will prove useful and it will show that translation invariant valuations do satisfy some analog to Theorem 2.8 (see Theorem 3.9). We refer to [24, Ch. 4] for a study of continuous and translation invariant valuations on $\text{Pa}(\{e_{i}\}_{i=1}^{n})$.
@@
Theorem 3.6 can be interpreted as a version of McMullen’s Theorem on decomposition for valuations on polytopes (see [3]). We characterize the homogeneous parts of the decomposition as separately additive functions of different dimensions, specifically when working within $\text{Pa}(\{e_{i}\}_{i=1}^{n})$. This formulation serves as a substitute for McMullen’s Theorem and we will use Theorem 3.6 in the proof of Theorem 1.1. This also leads to the following automatic continuity result.
@@
where we used the fact that additive functions are $\mathbb{Q}$-linear [26, Theorem 5.2.1]
@@ 4. A new characterization of the volume in $\mathbb{R}^{n}$
where $S(K)$ denotes the surface area of $K$ (see [40, Chapter 4]). On the other hand, since $B(0,r)\subset K\subset B(0,R)$, for $\varepsilon$ small enough we get
@@
(see [40, Lemma 2.3.6]). Combining these results we get
@@
Again, we restate that in the previous Lemma 4.3 the valuation condition can be relaxed, we only need the mapping $V$ to be weakly additive. This property is strictly weaker than being a valuation (see [35, Page 173] for an example of a weakly additive function on $\mathcal{K}^{n}$ which is not a valuation). Nevertheless, if $V$ is weakly additive, it is straightforward to see that the restriction $V|_{\text{Pa}(\{e_{i}\}_{i=1}^{n})}$ is a valuation on $\text{Pa}(\{e_{i}\}_{i=1}^{n})$, allowing us to apply the results from Section 3
@@ 5. Degrees of homogeneity for a bounded valuation
To see that this set cannot be reduced any further, we construct bounded and translation invariant valuations, $\phi_{l}$, that are $m$-homogeneous for any $m\in[0,n-1]$. These examples show both that each degree in this interval actually occurs, and that the argument from Lemma 4.3 can not be pushed to extend the values of $m$. Indeed, each $\phi_{l}$ vanishes on polytopes while satisfying $\phi_{l}(B)>0$. These valuations are analogous to the $p$-affine surface areas (see [28, Section 3.3] for a gentle introduction) and are based on the Gauss Kronecker curvature. Such valuations have been previously introduced in [33] and, in the case $n=2$, have been characterized as the upper semicontinuous valuations that are rotation invariant and homogeneous of degree different from 0, 1 or 2 (see [32]).
@@
A well known result due to Aleksandrov (see [4] or [40, Theorem 2.6.1]) states that the boundary of a convex body is twice differentiable almost everywhere with respect to the (n1)(n-1)-dimensional Hausdorff measure. This implies that the curvature is well defined almost everywhere, and moreover, $x\mapsto\kappa(K,x)$ is measurable. This mapping can also be described as the density of the absolutely continuous part of the curvature measure $C_{0}(K,\cdot)$ with respect to the (n1)(n-1)-dimensional Hausdorff measure. We refer the reader to [40, Chapter 4] and [21, Theorem 3.2] for the definition of $C_{0}(K,\cdot)$ and details. As a consequence, we have the following.
@@
where $\chi_{A}(x)$ denotes the indicator function of the set $A$. Applying the property [7, Theorem 3.6.1] we get that
@@
This is well defined since $\kappa(K,\cdot)$ is measurable and nonnegative, also, in the case $l=0$, we adopt the convention that $0^{0}=0$. We will need the following lemma taken directly from [41, Lemma 5].
@@
With this lemma and following the same procedure as in [41] we can prove that the previous mapping is a valuation.
@@ 6. Some negative results
The proof is based on [36, Theorem 1], with minor modifications, we include it here for completeness.
@@
Let $u=\sum_{i=1}^{n}\eta_{i}e_{i}\in\mathbb{S}^{n-1}$ with $\eta_{i}\neq 0$ for all $i=1,\dots,n$. We take $e_{i}^{\prime}=-\mathrm{sign}(\eta_{i})e_{i}$ and a simplex $T$ with $u,e_{1}^{\prime},\dots,e_{n}^{\prime}$ as its normal vectors. By [40, Lemma 5.1.1], taking $F_{i}=F(T,e_{i}^{\prime})$ and $F_{0}=F(T,u)$, we have
@@
Similarly, the mapping $\mu\mapsto\mu^{s}$, which gives the singular part of any measure with respect to the Haar measure, is linear. This follows from the uniqueness of the Lebesgue decomposition [9, Theorem 4.3.2]. Finally, the last mapping is defined as
@@
Besides Minkowski addition, Blaschke addition is another semigroup operation inside $\mathcal{K}^{n}$, that has proven to be quite useful. For the definition we refer the reader to [40] and [36]. We will see that a valuation that is Blaschke additive and bounded does not have to be continuous. To exhibit such an example, we will make use of McMullen’s Characterization of (n1)(n-1)-homogeneous continuous and translation invariant valuations, Theorem 2.7. The idea is to build valuations via the surface area measure, $S_{n-1}$, which is Blaschke additive, and then modify the function in the integral representation to obtain a valuation that is Borel measurable and bounded but not continuous.
@@
for any $K\in\mathcal{K}^{n}$. Since any balanced measure can be written as the difference of two surface area measures (check [36, Theorem 3]) we get that
@@
Now, let $L=\{\langle u,\cdot\rangle,u\in\mathbb{R}^{n}\}\subset C(\mathbb{S}^{n-1})$ be the subspace of linear functions. Note that this has finite dimension, which means that it is closed in all vector topologies. Now, the polar of $L$ is exactly the set of balanced measures and fgf-g is in the polar of balanced measures. By the Bipolar Theorem (see [12, Theorem 3.38]) we get that fgf-g is in $L$, that is, fgf-g is continuous and, since $g$ was continuous, we get that $f$ is continuous.
The converse result is immediate from the fact (see [40, Theorem 4.2.1]) that convergence of convex bodies $K_{j}\xrightarrow[]{}K$ in the Hausdorff metric implies that, for every continuous $f:\mathbb{S}^{n-1}\to\mathbb{R}$,
@@
Therefore, $\phi_{\chi_{\bigcup_{j\in\mathbb{N}}A_{j}}}$ is the pointwise limit of measurable functions and thus measurable. These facts show that $\mathcal{D}$ is a Dynkin class that contains all open sets. By a classical result of Dynkin classes (check [9, Chapter 1.6]), all Borel sets belong to $\mathcal{D}$.